What is the pH of of a solution of 0.60 M ammonium chloride ?


1
Expert's answer
2021-04-25T06:15:29-0400

Solution:

NH4Cl ⇌ NH4+ + Cl

1) NH4+ + H2O ⇌ NH3 + H3O+, KH = ???




2) NH3 + H2O ⇌ NH4+ + OH, KB = 1.8×10−5



3) H2O ⇌ H+ + OH, KW = 1.0×10−14





Hence,

KH = KW / KB = (1.0×10−14) / (1.8×10−5) = 5.56×10−10


NH4+ + H2O ⇌ NH3 + H3O+

From this, we can construct an ICE(Initial, Change, Equilibrium) Table:



From the ICE Table,

[NH3] = [H3O+] = x

[NH4+] = 0.60 - x


KH = [NH3] × [H3O+] / [NH4+]

5.56×10−10 = x2 / (0.60 - x)

We can assume that (0.60 - x) = 0.60

5.56×10−10 = x2 / 0.60

x2 = 3.336×10−10

x = 1.827×10−5


Hence,

x = [H3O+] = 1.827×10−5


We can convert between [H3O+] and pH using the following equation:

pH = - log[H3O+]

pH = - log(1.827×10−5) = 4.74

pH = 4.74


Answer: pH = 4.74